OpenStudy (anonymous):

Q: A 25.5 mL aliquot of HCl (aq) of unknown concentration was titrated with 0.113 M NaOH (aq). It took 51.2 mL of the base to reach the endpoint of the titration. The concentration (M) of the acid was __________. A: 0.227 Please help explain how to get the answer.

OpenStudy (anonymous):

Take 0.113M * 51.2ml = (M1) * (25.5) Then Divide (0.113*52.2) by 25.5 which = .2268 which = 0.227M

OpenStudy (anonymous):

Thanks you have been really helpful!

OpenStudy (anonymous):

your welcome no worries